OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18, ...

OpenStudy (zzr0ck3r):

-10/60=-1/6 (5/3)/-10 = (5/3)*(-1/10) = -1/6 (-5/18)*(3/5) = -1/6 so common ratio = |-1/6|=1/6 < 1 so we have that the series is converging because its geometric with ratio r st |r|<1 so its sum \[\frac{60}{1-(\frac{-1}{6}}\]

OpenStudy (zzr0ck3r):

\[\frac{60}{\frac{7}{6}}=60*\frac{6}{7}\]

OpenStudy (anonymous):

Thanks! Quick question, for the limit the formula is a1/1-r right?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

yes, for the sum...not the limit

OpenStudy (anonymous):

I need to find the limit, how would I do that?

OpenStudy (anonymous):

nvm I got it